Solve for x and y:
(x)^2 / 6 - y / 4 = 1,
x + y = 5

Answer by edjones(8007)

(x)^2 / 6 - y / 4 = 1,
x + y = 5
.
y=5-x
x^2/6 - (5-x)/4 = 1
2x^2-3(5-x)=12 Multiply each side by 12.
2x^2+3x-15=12
2x^2+3x-27=0
2*-27=-54; Two factors of -54 whose sum is 3 are 9 and -6
2x^2-6x+9x-27=0
2x(x-3)+9(x-3)=0
(2x+9)(x-3)=0
x=-9/2, y=19/2
x=3, y=2
.
Ed